• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Simonetti

    Log entry time 15:55:15 on November22,2008

    Entry number 250343

    keyword=Day shift summary

    Day shift summary

    08:00 Production continues
    10:00 Call the page number of "Tech on call" Todd Ewing about the vacuum
    10:35 Todd Ewing call: he will come at 12:00 to check the vacuum problem, and we will ask to MCC a controlled access, before he comes
    11:45 Stop run 3816 to pass in controlled access status and check the spectrometer vacuum problem
    12:13 Vacuum problem solved: a valve between the turbo pump and spectrometer malfunctioned in the closed position, Todd replaced the valve and the vacuum has returned to a reading of 0.
    12:16 Target moved to BeO target to tune the beam
    12:25 Target moved to He3 target to chek beam position
    12:35 Start production run 3817
    15:14 Alexandre request no beam to perform a background test; stop run 3819
    15:18 Background test completed; restart the production from run 3820